OpenStudy (marcon):

Sandra exclaims that her quadratic with a discriminant of –4 has no real solutions. Sandra then puts down her pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Sandra, in calm and complete sentences, how to find the solutions, even though they are not real. @Soniatonia

OpenStudy (anonymous):

Ok, so In the equation x² +2x + 2 = 0 the discriminant equals b² - 4*a*c = -4, that is because the square root requires a negative number

OpenStudy (anonymous):

the solution will not be real.

OpenStudy (anonymous):

because there is a solution already

OpenStudy (marcon):

How did you get x^2+2x+2=0 ?

OpenStudy (anonymous):

In math 'i' is the square root of minus 1

OpenStudy (anonymous):

The equation is an example of a quadratic with -4 discriminant, which you can use to to find the unreal roots

OpenStudy (marcon):

Can we use another equation, for example x^2 + 4x + 4 = 0 Would this ^^ one work?

OpenStudy (anonymous):

it's says that help Sandra find the solutions, even though they are not real, that's how I made x² +2x + 2 = 0 because of b² - 4*a*c = -4

OpenStudy (anonymous):

and yes we can

OpenStudy (anonymous):

But make sure the discriminant has to equal -4

OpenStudy (marcon):

so it would be x^2 + 4x + 4 = -4 ???

OpenStudy (anonymous):

i² = -1.

OpenStudy (marcon):

ok so what would the full answer be?

OpenStudy (anonymous):

 the solutions are 2 + i and 2 -i